OpenStudy (anonymous):

Please help! Use the method of completing the square to Transform the quadratic equation into the equation form (x – p)2 = q. -9 + 12x - 6x^2 = 0 A) (x - 1)2 = -0.5 B) (x - 1)2 = -1.5 C) (x - 1)2 = 0.5 D) (x - 1)2 = 1.5

OpenStudy (dtan5457):

First. -9+12x-6x^2=0 Divide by the value of a (-6) 9/6-2x+x^2 Reorder x^2-2x+9/6 Complete the square. Get a new value of c using (b/2)^2 x^2-2x=-9/6 new value of c=1, add to both sides x^2-2x+1=-9/6+1 x^2-2x+1=-1/2 Factor the trinomial (x-1)^2=-1/2

OpenStudy (anonymous):

Oh I see thank you!

OpenStudy (wolf1728):

Let's first transpose the equation into -6x^2 + 12x -9 = 0 To complete the square http://www.1728.org/quadr2.htm Move non-x term to the right -6x^2 + 12x = +9 Divide the equation by the coefficient of x^2 x^2 -2x = -1.5 take the coefficient of x divide it by 2 square it add it to both sides of the equation coefficient of x is -2 divided by 2 is -1 -1 squared is +1 so x^2 -2x +1= -.5

OpenStudy (wolf1728):

(x-1)*(x-1) = -.5 which is what dtan5457 calculated great minds think alike!

OpenStudy (anonymous):

Hahaha thanks wolf.